# SLoRD: Structural Low-Rank Descriptors for Shape Consistency in Vertebrae Segmentation

### v1

@@ 2.3 Contour Generation with Contour Priors
To precisely restore the contour, it is significant to accurately achieve the coefficient according to Eq (4). Besides, the contour generation requires the precise regression of spherical centers as described in Section 2.2. Thus, we devise the SLoRD framework by introducing the center decoder and coefficient decoder as shown in Figure 3(b). The ground truth of centers $C_{s}$ can be attained by Eq (6). And we employ the $L_{2}$ norm to regularize regressed centers $\hat{C}$, with the center loss $\mathcal{L}_{center}$ equal to $||\,C_{s}-\hat{C}\,||^{2}$.
@@
The direct loss regularization for combinational coefficients like the previous work [24] cannot work in this task. We reckon that phenomenon results from the magnitude difference between different dimensions of coefficient vectors and the magnitude variance between combinational coefficients from the same dimension. Thus, we transform coefficients into the radial distance with data-driven contour priors, then into the contour point cloud [25] centered at spherical centroids. After that, an explicit contour constraint $\mathcal{L}_{contour}$ is devised based on the average minimal distance from the contour point set $\mathcal{P}_{c}$ to the boundary interface $\mathcal{B}$ as illustrated by the following formula:
@@
Where $|\mathcal{P}_{c}|$ means the voxel number of $\mathcal{P}_{c}$. Besides, it might be confusing why we do not adopt a direct distance regularization on the regressed point clouds compared with ground truths. A comprehensible perspective is that the SLoRD network do not bear any perception ability to identify the sampled angle grid in the spherical coordinate system since no relevant angular priors are introduced. That is why basic contour descriptors encoded by the spherical coordinate system will boost the contour representation of vertebral anatomies.
To reduce computational costs, we uniformly sample $\tfrac{1}{3}$ boundary points from $\mathcal{B}$ to calculate the contour loss. Besides, an additional binary mask supervision $\mathcal{L}_{mask}$ (cross-entropy loss plus Dice loss) will enforce an implicit regularization to the coefficient decoder by introducing vertebral shape information. Thus, the total loss is defined as follows:
@@
Furthermore, segmentation inconsistency tends to appear in a sequential of vertebrae as highlighted in Figure 1. Some poor predictions even depict wrong identifications for a series of vertebrae. Under this circumstance, the iterative refinement for one vertebra each time cannot address this type of failure cases, and that is why we attempt to handle three vertebrae with sequential labels once. Assuming that coarse predictions reveal satisfactory performance for the top two vertebrae, even with a moderate inconsistency, SLoRD can significantly boost segmentation results.
@@ 2.4 Inference
In the inference phase, the patch to be refined in SLoRD is cropped along each vertebral center from coarse predictions as shown in Figure 3(a). The SLoRD framework will first refine three spherical centers corresponding to sequential vertebrae according to the prediction of the center decoder. Then the coefficient branch will output $3\times N$ (N refers to the dimension of $\bm{\rho}$) contour points surrounding three vertebral boundaries with different labels as revealed in Figure 3(c). After that, Marching Cube [26] is implemented to generate the triangle mesh, which is further voxelized into a filled volume [27]. After that, we perform the binarization on the coarse mask and multiply it with the filled volume to realize a refined mask with good label consistency and shape restoration via the auxiliary connected component analysis.
Table 1: Comparisons with other single-stage, two-stage and multi-stage methods on the public and hidden test datasets of VerSe 2019 (A-Dice: average Dice. M-Dice: median Dice. A-HD: average Hausdorff distance. M-HD: median Hausdorff distance).
@@
Two-stage Ours 90.94 95.08 6.14 4.74 90.87 95.90 6.53 4.25
@@ 3.1 Experimental Settings
Dataset. To evaluate the performance of our network, we conduct experiments on the challenging dataset [18]: MICCAI Challenge VerSe 2019, containing different pathological images. Specifically, VerSe 2019 consists of 160 spinal CT scans, which are split into 80/40/40 scans for training, public-test, and hidden-test cases in the first stage of generating coarse predictions. To ensure relatively fixed FOVs of iterative cropped patches in stage 2, we resample all data with a voxel space equal to $1.0\times 1.0\times 1.998mm^{3}$ in z, y, and x orientations. Then we select a size of $112\times 128\times 96$ to crop patches containing three sequential labeled vertebrae. And each patch is centered at the spherical centroid of the intermediate vertebra. To generate value variance in the center regression decoder, we exert a random translation on the patch center, with a range of $[-5,5]$ on each coordinate dimension. And there are 113 cervical, 436 thoracic, 314 lumbar vertebrae respectively. For each middle vertebra, we repeat the translation process 3 times. Thus, there re are a total of 2589 patches, which are split as 2100/489 cases for training and validation. During inference, the second-stage network is adopted to refine coarse predictions. Moreover, we choose the Dice score and Hausdorff distance (HD) as quantitative metrics [29].
@@
Implementation Details. In stage 1, we choose MedNeXt as the baseline model, trained for 1000 epochs with $10.53M$ parameters. And we utilize a combination of cross entropy loss and Dice loss followed by [20]. The patch size is $128\times 160\times 96$. In stage 2, the refinement network is trained for 400 epochs, which employs the encoder of 3D UNet with only $5.15M$ parameters. And we utilize a weighted loss $\mathcal{L}_{total}$ as illustrated in Section 2.3. We train all models using AdamW optimizer. With the linear warm-up strategy, the initial learning rate is set as $-45e\mbox{-}$ 4 with a cosine learning rate decay scheduler, and weight decay is set as $-51e\mbox{-}$ 5. The batch size is 2 and 16 for two stages individually. All models are implemented based on Pytorch and trained on 2 NVIDIA Tesla V100 GPUs.
Figure 4: Qualitative visualizations of models in the benchmark.
Figure 6: Visualizations of the decomposed contour descriptors. Different indexes refer to different contour vectors.
3.2 Experimental Results
@@
To evaluate the efficacy of our proposed method, we conduct qualitative and quantitative experimental results on VerSe 2019. As illustrated by Table 1, our method achieves SOTA performance on the average Dice and Hausdorff distance (HD) of the hidden test dataset compared with other single-stage and multi-stage methods. Specifically, our framework surpasses MedNeXt [22] (the best single-stage model) by $1.96\%\uparrow$ average Dice score and $4.82mm\downarrow$ average Hausdorff distance. And these two metrics come to $0.93\%\uparrow$ and $0.43mm\downarrow$ when compared with Spine-Transformer [15] (the best two-stage model), $8.31\%\uparrow$ and $3.45mm\downarrow$ compared with Chen’s method [18] (the best multi-stage model). Moreover, the median Dice and HD of our method also reach a promising performance. Obviously, the superiority of segmentation performance originates from the consistency regularization generated by our instance-based SLoRD architecture. By labeling sequential instance masks, we will acquire semantic segmentation masks with better consistency.
Qualitative visualizations are also displayed in Figure 1 and 4, which reveals that our framework could attain more consistent predictions in each vertebra even for challenging cases. As revealed in Figure 5, the whole refinement process is visualized. The results in the dashed box represent the efficacy of our proposed SLoRD. It is worth highlighting that each instance mask is generated via the linear combination of basic contour descriptors, which can be achieved in a data-driven manner. Here we give a depiction of decomposed contour priors by SVD as shown in Figure 6.
Besides, the efficiency analysis is conducted between our devised pipeline and other methods in the VerSe 2019 challenge. Specifically, we choose network parameters (M), computational costs (G FLOPs), and inference time per case (s) as quantitative metrics. As illustrated by Table 2, our approach only requires $25.18$M network parameters, then achieves a $142.8$s average inference time for each CT scan, which is much faster than Payer’s method (260.0s) and Tao’s method (176.3s).
Table 2: Left: Ablation study on the plug-and-play property of SLoRD when applied to other CNNs and Transformer-based networks. Quantitative results on the hidden test dataset are provided. Right: The efficiency evaluation between our method and other multiple-stage approaches on the hidden test datasets of VerSe 2019.
Model A-Dice M-Dice A-HD M-HD
nnUNet [20] 86.59 92.62 12.78 12.00
nnUNet [20] + SLoRD 88.85 93.31 8.48 6.39
Verteformer [7] 86.54 90.74 10.55 10.51
Verteformer [7] + SLoRD 89.13 94.26 7.46 7.23
@@
Table 3: Left: ablation study on the orientation of positive Z-axis, sampling interval, and the number of basic contour descriptors. Right: ablation study on the type of centers between the centroid and spherical center (ASD [29]: average surface distance).
@@
Figure 8: Improvement of our proposed SLoRD when applied to coarse predictions from other segmentation networks including the single-stage Verteformer [7], two-stage Spine-Transformer [15], and multi-stage Payer’s methods [18].
3.3 Ablation Studies
@@
In Table 3, we investigate a detailed ablation analysis on the contour representations based on the spherical coordinate system. For a simple yet effective evaluation, we adopt the average surface distance (ASD) between contour points and the boundary to assess the contour restoration performance. Here the ASD metric is calculated via the linear combination of full-rank contour descriptors by SVD. Thus, this distance value stands for the best performance that can be theoretically achieved due to the non-convex shape property of vertebral anatomies.
@@
1) The orientation of the positive Z-axis: the voxel z-axis as the objective orientation achieves the lowest ASD value. Compared with the voxel y-axis, there is a $0.025$ voxel distance drop. Besides, as revealed in Figure 7, there will be a deviation for contour points adjacent to the left and right ends of vertebrae. 2) The sampling interval of the polar angle $\phi$ and the azimuth angle $\theta$: different values of the angular interval are considered. As shown in Table 3, a smaller interval means a denser sampling for the contour point cloud, which will inevitably promote the precision of regressed contours by SLoRD, and then elevate the final segmentation performance. Considering the computational efficiency, the hyper-parameter is specified as $5^{\circ}$, then the dimension $N$ of each contour descriptor $\rho$ is equal to $\tfrac{360^{\circ}}{5^{\circ}}\times(\tfrac{180^{\circ}}{5^{\circ}}+1)=2664$. 3) Spherical centroid vs. Centroid: adopting the spherical centroid reduces the ASD value [29] in low-rank and full-rank approximations, with a $0.053$ and $0.267$ decrease respectively. That demonstrates the advancement of spherical centroids owing to the optimization objective of its design.

### v2

@@ 3.3 Contour Generation with Explicit Contour Priors
To precisely restore the contour with explicit contour descriptors, it is significant to accurately achieve the coefficient according to Eq (3). Besides, the contour generation requires the precise regression of spherical centers as described in Section 3.2. Thus, we devise the SLoRD framework by introducing the center decoder and coefficient decoder as shown in Fig.4(c).
Center decoder: Deep features from the encoder are transformed into regressed spherical centers via combinative operations of convolution, batch normalization, average pooling. The pooling process projects the feature dimension as 3, representing coordinates of the regressed center $\hat{C}=(\hat{c}_{x},\hat{c}_{y},\hat{c}_{z})$. The ground truth of center $C_{s}$ can be attained by Eq (4). And we employ the $L_{2}$ norm to regularize regressed centers $\hat{C}$, with the center loss $\mathcal{L}_{center}$ equal to $||\,C_{s}-\hat{C}\,||^{2}$. Concerning the limitation that some methods wrongly detect centers of specific pathological vertebrae, we aim to introduce Gaussian positional prompts as additional inputs for positional guidance. As shown in Fig.4(b), the positional prompt is defined as a Gaussian distribution $\mathcal{G}$ based on the follow formula:
$\mathcal{G}(\bm{\mu};\bm{\sigma})=\mathcal{G}(\mu_{x};\sigma_{x})\cdot\mathcal{G}(\mu_{y};\sigma_{y})\cdot\mathcal{G}(\mu_{z};\sigma_{z})$ (5)
[…]
Specifically, the mean value $\bm{\mu}=(\mu_{x},\mu_{y},\mu_{z})$ is set as the centroid based on the coarse prediction. For defining the variance $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$, we aim to cover the whole shape distribution using the given $\mathcal{G}$. Thus, as suggested by the two-sigma rule in mathematical statistics, the length of $4\bm{\sigma}$ should encompass the 3-dimension length $\bm{m}=(m_{x},m_{y},m_{z})$ of each instance mask. Further, a threshold $\bm{\bar{m}}$ is introduced to ensure a wide enough region with attention weights decaying from the center $\bm{\mu}$ to the vertebral boundary. The center decoder will modify coarse centroids $\bm{\mu}$ as refined spherical centroids via $\mathcal{L}_{center}$.
@@
Coefficient decoder: Similar to the center decoder, this part is also a combination of convolution, normalization and average pooling layers, with the projected dimension equal to the rank size. Here a direct loss on combinational coefficients like the previous work [30] cannot work in this task. We reckon that phenomenon results from the magnitude difference between different dimensions of coefficient vectors and the magnitude variance between coefficient vectors at the same dimension. Thus, we transform coefficients into the radial distance with data-driven contour priors, then into the contour point cloud [41] centered at spherical centroids. After that, an explicit contour constraint $\mathcal{L}_{contour}$ is devised based on the average minimal distance from the contour point set $\mathcal{P}_{c}$ to the boundary interface $\mathcal{B}$ as illustrated by the following formula:
@@
where $|\mathcal{P}_{c}|$ means the voxel number of $\mathcal{P}_{c}$. Besides, it might be confusing why a straightforward distance regularization is not conducted on the regressed point clouds. A comprehensible perspective is that SLoRD does not bear any perception ability to identify the sampled angle grid in the spherical coordinate system with no angular priors introduced. That is why basic contour descriptors will boost the contour representation of vertebral anatomies.
Loss term: To reduce computational costs, we uniformly sample $\tfrac{1}{3}$ boundary points from $\mathcal{B}$ to calculate the contour loss. Besides, an additional binary mask supervision $\mathcal{L}_{mask}$ (cross-entropy + Dice loss) will exert implicit shape information. Thus, the total loss is defined as follows:
@@
Furthermore, segmentation inconsistency tends to appear in a sequence of vertebrae as highlighted in Fig.1. Under this circumstance, the iterative refinement for one vertebra each time cannot well address this type of failure cases, hence we attempt to handle three sequential vertebrae once. Assuming that coarse predictions reveal satisfactory performance for the top two vertebrae, even with a moderate inconsistency, SLoRD can significantly boost segmentation results.
Figure 5: Iterative refinement for coarse masks in the inference stage.
@@ 3.4 Inference
In the inference phase, the patch to be refined in SLoRD is cropped along each vertebral center from coarse predictions as shown in Fig.4(a). The SLoRD framework will first refine three spherical centers corresponding to sequential vertebrae according to the prediction of the center decoder. Then the coefficient branch will output $3\times N$ (N refers to the dimension of $\bm{\rho}$) contour points surrounding three vertebral surfaces with different labels as revealed in Fig.5. After that, Marching Cube [25] is implemented to generate triangle meshes, which are further voxelized into a filled volume [33]. After that, we perform the binarization on the coarse mask and multiply it with the filled volume to realize a refined mask with good label consistency and shape restoration via the auxiliary connected component analysis.
Table 1: Comparisons with other single-stage, two-stage, and multi-stage methods on the public and hidden test datasets of VerSe 2019 (A-Dice: average Dice. M-Dice: median Dice. A-HD: average Hausdorff distance. M-HD: median Hausdorff distance).
@@
Two-stage Ours 92.17 95.62 5.91 4.06 91.33 96.03 5.97 4.27
Figure 6: Qualitative visualizations of hard cases by single-stage, two-stage, and multi-stage models in the benchmark. The $1^{st}$ row: a case from VerSe 2019. The $2^{nd}$ and $3^{rd}$ row: two cases from VerSe 2020.
@@ 4.1 Experimental Settings
VerSe 2019 & 2020. To assess the efficacy of our method, we conduct experiments on the Large Scale Vertebrae Segmentation Challenge (VerSe 2019 & 2020) [34]. Specifically, VerSe 2019 consists of 160 spinal CT scans, split into 80/40/40 scans for training, public test, and hidden test cases. VerSe 2020 is a more challenging dataset, containing more CT scans with pathological appearances, fractures, and metal artifacts. We follow the standard data-split setting, with 89/103/103 scans for training, public test, and hidden test cases.
Table 2: Comparisons with other single-stage, two-stage, and multi-stage methods on the public and hidden test datasets of VerSe 2020 (A-Dice: average Dice. M-Dice: median Dice. A-HD: average Hausdorff distance. M-HD: median Hausdorff distance).
Stage Method Public test Hidden test
3D UNet [11] 82.69 89.49 11.15 9.40 83.32 88.90 10.59 9.02
Xiangshang Z. [34] 83.58 92.69 15.19 9.76 85.07 93.29 12.99 8.44
nnUNet [15] 85.34 92.10 10.01 8.47 86.83 91.79 9.29 7.58
Single-stage Verteformer [45] 85.44 92.04 9.19 8.86 86.57 91.47 9.80 8.23
UNeXt[39] 81.19 90.30 13.39 11.94 82.77 87.94 12.39 12.72
Swin UNETR [37] 82.57 90.68 12.07 10.16 81.83 88.79 12.98 11.37
MedNeXt [32] 87.35 93.26 8.98 7.50 87.75 92.41 8.31 6.28
Multi-stage Sekuboyina A. [34] 78.05 85.09 10.99 6.38 79.52 85.49 11.61 7.76
Two-stage Yeah T. [34] 88.88 92.93 9.57 5.43 87.91 92.76 8.41 5.91
Multi-stage Payer C. [31] 91.65 95.72 5.80 4.06 89.71 95.65 6.06 3.94
Multi-stage Chen D. [34] 91.72 95.52 6.14 4.22 91.23 95.21 7.15 4.30
Two-stage Tao R. [38] 91.29 95.33 6.30 4.73 91.65 94.72 6.29 5.07
Two-stage Ours 92.09 95.75 5.30 3.78 92.31 95.33 5.55 3.74
@@
To ensure relatively fixed FOVs of iterative cropped patches in the second stage, we resample all data with a voxel space equal to $1\times 1\times 1.998mm^{3}$ and $0.8262\times 0.8262\times 1mm^{3}$ for VerSe 2019 and 2020 respectively. Then we crop patches containing at least three sequential vertebrae, with the patch size equal to $112\times 128\times 64$ and $128\times 128\times 96$. Each patch is centered at the spherical centroid of the intermediate vertebra. A random translation is exerted on the patch center to generate the variance in the center regression decoder, with a $[-5,5]$ range on each coordinate axis. For each middle vertebra, we repeat the translation process 3 times. To further simulate various types of imprecise boundary prediction, different pretrained weights are adopted to generate coarse predictions for the positional prompt. For VerSe 2019, a total of 7767 patches from 80 training CT scans are split as 6214/1553 cases for training and validation. For VerSe 2020, following the same $4\mbox{:}1$ ratio, 10188 patches are split as 8150/2038 cases. During inference, the SLoRD framework is adopted to refine coarse predictions. Moreover, Dice score and Hausdorff distance (HD) are chosen as quantitative metrics [44].
@@
Implementation Details. All experiments are implemented based on Pytorch and trained on 2 NVIDIA Tesla V100 GPUs. In stage 1, MedNeXt [32] is selected as the U-shape backbone with large receptive fields to extract long-term semantics, trained for 1000 epochs with $10.53M$ parameters. And we utilize a combination of cross entropy loss and Dice loss followed by [15]. The patch size is $128\times 160\times 96$. In stage 2, SLoRD is trained for 400 epochs, which employs the encoder of 3D UNet with only $5.15M$ parameters. And we utilize a weighted loss $\mathcal{L}_{total}$ as illustrated in Section 3.3. We train all models using AdamW optimizer. With the linear warm-up strategy, the initial learning rate is set as $-45e\mbox{-}$ 4 with a cosine learning rate decay scheduler, and weight decay is set as $-51e\mbox{-}$ 5. The batch size is 2 and 16 for two stages individually. The hyperparameters $s$ (sampling interval) and $k$ (number of basic contour descriptors) are set as $5^{\circ}$ and $200$ for a better trade-off between performance and efficiency.
4.2 Performance Comparison
4.2.1 VerSe 2019
@@
To evaluate the efficacy of SLoRD, quantitative experimental results are first conducted on VerSe 2019. As illustrated by Table 1, our method achieves SOTA performance on the average Dice and Hausdorff distance (HD) of the hidden test dataset compared with other single-stage, two-stage, and multi-stage methods. Specifically, our framework surpasses MedNeXt [32] (the best single-stage model) by $2.42\%\uparrow$ average Dice score and $5.38mm\downarrow$ average Hausdorff distance. And these two metrics come to $1.39\%\uparrow$ and $0.99mm\downarrow$ when compared with Spine-Transformer [38] (the strongest two-stage model), $1.53\%\uparrow$ and $1.11mm\downarrow$ compared with Payer’s method [34] (the powerful multi-stage model). Moreover, the median Dice and HD of our method also reach a quite promising performance.
4.2.2 VerSe 2020
To further validate the robustness and universality of SLoRD, we carry out a quantitative analysis on VerSe 2020, which contains more challenging cases with deformed shapes and metal artifacts. As shown in Table 2, our approach achieves the best boundary delineation results with the lowest average HD values. Compared with Chen’s method [34], $0.84mm\downarrow$ and $1.60mm\downarrow$ Hausdorff distance are acquired for the public and hidden test datasets. Also, the median HD values reveal our method could improve the segmentation performance for not only satisfactory cases but poor cases. Obviously, the superior segmentation performance originates from the strict voxel consistency and contour precision regularization generated from the instance-based SLoRD architecture. By labeling sequential instance masks, we will acquire semantic masks with better consistency.
Table 3: The model efficiency evaluation between our method and other approaches on the hidden test datasets of VerSe 2019.
@@ 4.2.2 VerSe 2020
4.2.3 Qualitative results
Qualitative visualizations of some challenging cases are displayed in Fig.6. The first case visualizes thoracic vertebrae with similar shapes. Besides, the spinous process of the precedent vertebra shows the same height as the superior articular process and vertebral body of the latter vertebra, prone to imprecise boundary predictions. The second case is plagued by metal artifacts, making it difficult to identify and locate several vertebrae in the bottom. The third case illustrates the potential problem of inter-class boundary confusion between normal vertebrae and transitional vertebrae [34]. And our method can better address the above obstacle with intra-vertebrae segmentation consistency.
For a detailed mask refinement process, intermediate results are provided in Fig.7. Given coarse predictions in stage 1, SLoRD will generate contour predictions for each vertebra. Then discrete contour point clouds are transformed into a sequence of masked meshes. After filling the inner volume, refined masks will be attained with good boundary precision for each class as shown in the dashed box of Fig.7. SLoRD can function well as long as the top two vertebrae perform satisfactorily, even with moderate inconsistency. It is worth highlighting that each instance mask is generated via the linear combination of basic contour descriptors, which can be acquired by SVD offline. A depiction of decomposed contour priors can be referred to in Fig.3.
4.2.4 Model Efficiency
Besides, the efficiency analysis is conducted between our framework and other typical methods on the hidden test dataset of VerSe 2019. Specifically, we choose model parameters, computational costs, and inference time per case as quantitative metrics. As illustrated by Table 3, our approach only requires $25.18$M network parameters, which is the most lightweight model in the referenced methods. Besides, our approach achieves a $142.8$s average inference time for each CT scan, much faster than Payer’s method ($260.0$s) and Tao’s method ($176.3$s).
Table 4: Ablation study on the orientation of positive Z-axis, sampling interval, and the number of basic contour descriptors (ASD [44]: average surface distance).
@@ 4.2.4 Model Efficiency
4.3 Discussion
4.3.1 Settings for the spherical-based contour representations
@@
We investigate a detailed ablation analysis on the contour representations based on the spherical coordinate system. For a simple yet effective evaluation, we adopt the average surface distance (ASD) between contour points and the boundary to assess the contour restoration performance. The ASD value in Table 4 is calculated via the linear combination of full-rank contour descriptors from SVD. Thus, this distance value stands for the best performance that can be theoretically achieved due to vertebral non-convex shape property and deviation due to integer approximation.
The orientation of the positive Z-axis in the spherical coordinate system: As revealed in Table 4, the voxel z-axis as the objective orientation achieves the lowest ASD value. Compared with the voxel x-axis, there is a $0.227$ voxel distance drop. Besides, as revealed in Fig.8, there will be a deviation for contour points adjacent to the spinous process of vertebrae when comparing the setting of z-axis with y-axis.
@@
Sampling interval of the azimuth angle $\phi$ and the polar angle $\theta$: Different values of the angular interval are also considered. As shown in Table 4, a smaller interval means a denser sampling for the contour point cloud, which will promote the precision of regressed contours by SLoRD, and then elevate the final segmentation performance. However, that behavior will inevitably bring heavier computational costs. Considering the balance, that hyper-parameter is specified as $5^{\circ}$, then the dimension $N$ of each contour descriptor $\bm{\rho}$ is equal to $\tfrac{360^{\circ}}{5^{\circ}}\times(\tfrac{180^{\circ}}{5^{\circ}}+1)=2664$.
The rank of basic contour descriptors: According to Table 4, 200 basic contour descriptors can already restore the original contour with sufficient precision. That is also proven in the qualitative results of Fig.3.
Spherical centroid vs. Centroid: Adopting the spherical centroid reduces the ASD value in both low-rank and full-rank approximations, with a $0.053\downarrow$ and $0.267\downarrow$ respectively as shown in Fig.9. That demonstrates the advancement of spherical centroids owing to the distance-based optimization objective between contour points and the vertebral surface.
Figure 9: Left: ablation study on the type of centers between the centroid and spherical center (ASD [44]: average surface distance). Right: performance comparison between our method and DeepSSM on VerSe 2020.
Figure 10: Qualitative results for the ablation study on the center decoder and mask decoder of SLoRD.
Table 5: Ablation study on the combination of decoders on VerSe 2020.
Method Dice score (%) $\uparrow$ $HD$ (mm) $\downarrow$
Cervical Thoracic Lumbar Mean Median Cervical Thoracic Lumbar Mean Median
Ours 90.27 93.44 90.69 92.31 95.33 6.37 4.68 6.04 5.55 3.74
w/o mask decoder 87.82 90.32 88.38 90.13 94.47 8.00 6.87 7.07 7.29 5.86
w/o coefficient decoder - - - - - - - - - -
w/o center decoder 87.41 90.50 85.96 88.64 93.80 8.13 6.14 9.11 8.09 6.75
w/o positional prompts 89.58 89.83 89.62 91.22 94.85 6.86 7.09 6.19 6.63 5.03
Table 6: Ablation study on the plug-and-play property of SLoRD when applied to other single-stage, two-stage, and multi-stage methods on vertebrae segmentation. Quantitative results on the hidden test dataset of VerSe 2019 and 2020.
VerSe 2019
Methods A-Dice M-Dice A-HD M-HD
Verteformer [45] 86.54 90.74 10.55 10.51
Verteformer [45] + SLoRD 89.34 92.98 6.73 6.31
Payer C. [31] 89.80 95.47 7.08 4.45
Payer C. [31] + SLoRD 90.97 95.26 5.93 4.28
Tao R. [38] 89.94 94.01 6.96 4.73
Tao R. [38] + SLoRD 91.46 95.72 5.73 4.23
VerSe 2020
Verteformer [45] 86.57 91.47 9.80 8.23
Verteformer [45] + SLoRD 89.22 93.90 7.64 5.55
Payer C. [31] 89.71 95.65 6.06 3.94
Payer C. [31] + SLoRD 91.72 95.73 5.58 4.08
Tao R. [38] 91.65 94.72 6.29 5.07
Tao R. [38] + SLoRD 92.24 95.11 5.27 4.34
4.3.2 Structural ablation on SLoRD
Furthermore, we conduct the structural ablation analysis for SLoRD on VerSe 2020.
Coefficient decoder: Without the coefficient decoder, contours cannot be generated. Thus, that is the essential part. For the reason why the regression of contour points is not performed directly in this part, we reckon that models show deficient abilities to perceive and localize contour points with the specific angle. Therefore, basic contour descriptors with latent angular information serve as strong guidance to generate target contour points.
Mask decoder: Without it, SLoRD reveals degraded segmentation performance, with a $2.18\%\downarrow$ average Dice score according to Table 5. Carrying out the cause and effect analysis, this mask decoder will bring additional shape information, which boosts the contour restoration precision according to Eq (3). For more detailed evaluations, the missing of this component will affect the boundary delineation for cervical vertebrae as shown in Fig.10.
Center decoder: Since the coupling effects between spherical centers and coefficients, removing the center decoder will diminish the segmentation performance significantly, with $3.67\%\downarrow$ average Dice score and $3.07mm\uparrow$ average HD. Besides, as portrayed in Fig.10, there exists a missing detection for a specific vertebra, resulting in sequential identification errors. That phenomenon significantly weakens segmentation results on lumber vertebrae [8].
Positional prompts: The positional priors promote the regression for spherical centroids, further boost the precision of contour generation. This detection guidance is beneficial to segment the thoracic region with adjacent vertebrae.
4.3.3 Plug-and-play